OpenStudy (anonymous):

induction step for 1*2^1+2*2^2+3*2^3+...+k*2^k=(k-1)2^(k-1)+2?

OpenStudy (anonymous):

I plugged k-1 in already and got \[(k-1)2^{k-1} + 2 + (k+1) * 2^{k+1}=k*2^k+2\]

OpenStudy (anonymous):

I just can't seem to simplify the left side to make it the same as the right

OpenStudy (anonymous):

\[(k*2^k)/2 -((5*2^k)/2)+1\] this is what I ended up on the left side

OpenStudy (asnaseer):

1st show its true for k=1 then assume its true for, say, k=j then prove its also true for k=j+1

OpenStudy (anonymous):

I dont think it is true for k=2 (2-1)2^(2-1) +2=1*2+2=4 not right..

OpenStudy (anonymous):

you are correct it is not true

OpenStudy (anonymous):

than induction wont work

OpenStudy (anonymous):

yeah I wrote it wrong it's suppose to be (k-1)2^(k+1)+2

OpenStudy (anonymous):

1*2^1+2*2^2+3*2^3+...+k*2^k=(k-1)2^(k+1)+2 but I still can't get it thought I could

OpenStudy (anonymous):

give me a sec I will write it down.

OpenStudy (anonymous):

so plugging in k+1 I get \[(k-1)*2^{k+1}+2 +(k+1)*2^{k+2}=k*2^k+2\]

OpenStudy (anonymous):

oops the right part is k*2^(k+2)+2

OpenStudy (anonymous):

yes and thats the end of induction. You showed that for k+1 that formula stands.

OpenStudy (anonymous):

no because I couldn't show that the left side is equal to the right..

OpenStudy (anonymous):

how do I simply so that it works?

OpenStudy (anonymous):

simplify*

OpenStudy (anonymous):

1*2^1+2*2^2+3*2^3+...+k*2^k=(k-1)2^(k+1)+2 so for k+1 we need that the result should be (k+1-1)2^(k+1+1) +2 and this is what you did

OpenStudy (anonymous):

but how about the left side? the only reason why I got (k)2^(k+2)+2 because I plugged in k+1 on the right side but don't you need to also prove that the left side equals the right if you get what I'm trying to say